Independent random samples of professional football and basketball players gave the following information. Assume that the weight distributions are mound-shaped and symmetric.

Weights (in lb) of pro football players: x1; n1 = 21

249 261 254 251 244 276 240 265 257 252 282
256 250 264 270 275 245 275 253 265 270

Weights (in lb) of pro basketball players: x2; n2 = 19

203 200 220 210 192 215 222 216 228 207
225 208 195 191 207 196 182 193 201

(a) Use a calculator with mean and standard deviation keys to calculate x1, s1, x2, and s2. (Round your answers to one decimal place.)

x1 =
s1 =
x2 =
s2 =


(b) Let μ1 be the population mean for x1 and let μ2 be the population mean for x2. Find a 99% confidence interval for μ1μ2. (Round your answers to one decimal place.)

lower limit    
upper limit    


(c) Examine the confidence interval and explain what it means in the context of this problem. Does the interval consist of numbers that are all positive? all negative? of different signs? At the 99% level of confidence, do professional football players tend to have a higher population mean weight than professional basketball players?

Because the interval contains only negative numbers, we can say that professional football players have a lower mean weight than professional basketball players.

Because the interval contains both positive and negative numbers, we cannot say that professional football players have a higher mean weight than professional basketball players.   

Because the interval contains only positive numbers, we can say that professional football players have a higher mean weight than professional basketball players.


(d) Which distribution did you use? Why?

The standard normal distribution was used because σ1 and σ2 are unknown.

The Student's t-distribution was used because σ1 and σ2 are known.    

The Student's t-distribution was used because σ1 and σ2 are unknown.

The standard normal distribution was used because σ1 and σ2 are known.
Answer #1
Football ( X ) Σ ( Xi- X̅ )2 Basketball ( Y ) Σ ( Yi- Y̅ )2
249 114.49 203 7.84
261 1.69 200 33.64
254 32.49 220 201.64
251 75.69 210 17.64
244 246.49 192 190.44
276 265.69 215 84.64
240 388.09 222 262.44
265 28.09 216 104.04
257 7.29 228 492.84
252 59.29 207 1.44
282 497.29 225 368.64
256 13.69 208 4.84
250 94.09 195 116.64
264 18.49 191 219.04
270 106.09 207 1.44
275 234.09 196 96.04
245 216.09 182 566.44
275 234.09 193 163.84
253 44.89 201 23.04
265 28.09
270 106.09
Total 5454 2812.29 3911 2956.56

Mean X̅ = Σ Xi / n
X̅ = 5454 / 21 = 259.7
Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 2812.29 / 21 -1 ) = 11.9

Mean Y̅ = ΣYi / n
Y̅ = 3911 / 19 = 205.8
Sample Standard deviation SY = √ ( (Yi - Y̅ )2 / n - 1 )
SY = √ ( 2956.56 / 19 -1) = 12.8

x1 =259.7 s1 = 11.9   x2 =205.8   s2 =12.8

Part b)

Confidence interval :-

Critical value   t(α/2, DF) = t(0.01 /2, 36 ) = 2.7195 ( From t table )



DF = 36


Lower Limit =
Lower Limit = 43.2
Upper Limit =
Upper Limit = 64.5
99% Confidence interval is ( 43.2 , 64.5 )

Part c)

Because the interval contains only positive numbers, we can say that professional football players have a higher mean weight than professional basketball players.

Part d)

The Student's t-distribution was used because σ1 and σ2 are unknown.